### v2

# Toward Carnot efficient high output power heat engines using bubbly two-phase flow

Toward Carnot efficient high output power heat engines using bubbly two-phase flow
Dror Miron The Nancy and Stephen Grand Technion Energy Program, Technion – Israel Institute of Technology, Haifa 3200003, Israel Lava Energy LTD. These authors contributed equally. Correspondence: sdmiron@campus.technion.ac.il Yuval Neumann The Nancy and Stephen Grand Technion Energy Program, Technion – Israel Institute of Technology, Haifa 3200003, Israel Lava Energy LTD. These authors contributed equally. Joseph Cassell Lava Energy LTD. These authors contributed equally. Nir Feintuch Faculty of Mechanical Engineering, Technion – Israel Institute of Technology, Haifa 3200003, Israel Lava Energy LTD. Alexey Shinkarenko Lava Energy LTD. Carmel Rotschild The Nancy and Stephen Grand Technion Energy Program, Technion – Israel Institute of Technology, Haifa 3200003, Israel Faculty of Mechanical Engineering, Technion – Israel Institute of Technology, Haifa 3200003, Israel Lava Energy LTD. Correspondence: carmelr@technion.ac.il
SUMMARY
Thermodynamic gas power cycles achieving Carnot efficiency require isothermal expansion, which is associated with slow processes and results in negligible power output. This study proposes a practical method for rapid near-isothermal gas expansion, facilitating efficient heat engines without sacrificing power. The method involves bubble expansion in a heat transfer liquid, ensuring efficient and near-isothermal heat exchange. The mixture is accelerated through a converging-diverging nozzle, converting thermal energy into kinetic energy, thereby rotating a reaction turbine for electricity generation. Nozzle experiments with air and water yielded a polytropic index $<$1.052, enabling up to 71% more work extraction than adiabatic expansion. Simulations indicate that utilizing these nozzles for thrust generation enables decreasing heat transfer irreversibilities in the heat engine, consequently resulting in up to 22.6% higher power output than an ideal heat engine based on the organic Rankine cycle. This work paves the way for an efficient and high-power heat-to-power solution.
Graphical abstract
KEYWORDS
Isothermal expansion, Heat engine, Waste heat recovery, Carnot efficiency, Two-phase nozzle
INTRODUCTION
To limit global warming corresponding to international goals, global greenhouse gas emissions should be urgently reduced 1. One of the promising methods to limit greenhouse gas emissions is by improving energy efficiency. As 72% of global primary energy production is lost to waste heat 2, there is great potential for energy to be sourced from waste heat recovery methods. Much of this potential lies in industrial applications, where heat is dissipated as a by-product in tens to hundreds of kilowatts, for which available heat engines are inefficient and expensive.
Due to its enhanced thermodynamic performance, power systems based on the organic Rankine cycle (ORC) have been considered promising for waste heat 3, 4 and geothermal energy recovery systems 5. Therefore, improving ORC systems’ performance received research attention in recent years 6, 7, 8, 9, 10, 11. Specifically, it has been suggested to implement quasi-isothermal expansion as an alternative to adiabatic expansion since more work can be extracted 12, 7. For example, studies indicated a flooded-expansion method to achieve quasi-isothermal expansion 13, 14, 7. Such cycles involve flooding the expansion device (screw or scroll expander) with a liquid that is in thermal equilibrium with the working fluid (WF). The liquid acts as a heat reservoir for the expanding vapor, which allows it to maintain its temperature during the expansion process. The isothermal expansion also produces super-heated vapor at the expander outlet, facilitating internal regeneration normally unavailable in conventional ORC 14. Although promising, studies investigating flooding the screw expander found no change to the outlet temperature (i.e., isothermal expansion has not been achieved) and a decreased trend of the cycle efficiency when increasing the flooding ratio15, 7, 8. The scroll expander showed deteriorating performance when increasing the flooding ratio due to pressure drop in the scroll expander inlet 12, 16. Another challenge mentioned in this configuration is phase separation after the expander. Consequently, a practical approach has yet to be known for realizing quasi-isothermal expansion in organic cycles.
In contrast to expander flooding, studies in marine propulsion have proposed employing air bubble expansion within water to achieve isothermal expansion. This method significantly increases the contact area between the fluids and facilitates efficient heat transfer. Additionally, this method substantially elevates the liquid-to-gas mass ratio relative to the liquid flooding technique, promoting gas expansion within a heat reservoir. Integrating these effects in a converging nozzle has enhanced thrust by allowing more effective work extraction compared to single-phase air ramjets 17, 18, 19. To further augment thrust, it has been proposed that choked flow conditions be attained. As such, a converging-diverging nozzle has been studied and experimented with, demonstrating supersonic flow and thrust enhancement 20.
This study presents the development of an efficient, high-power heat engine based on a novel thermodynamic vapor cycle incorporating quasi-isothermal expansion of bubbles in a heat transfer liquid (HTL). The large bubbles’ surface area supports efficient heat transfer between the phases, and the substantial mass and heat capacity discrepancies between the phases facilitate minimal temperature drop, thereby implementing a quasi-isothermal process. The two-phase mixture flows through converging-diverging nozzles, converting the thermal energy to thrust, which rotates a reaction turbine. To this end, we combine experimental pressure measurements and computational fluid dynamic (CFD) analysis of a two-phase flow in a converging-diverging nozzle to extract an upper boundary for the polytropic index of air and quantify the work enhancement of the near-isothermal compared to adiabatic expansion. Building on this, we simulate the performance of a heat engine that integrates the nozzles as expanders to drive a reaction turbine for electricity generation from a constant-temperature heat source. The results highlight the potential of this design, as a polytropic index of air $n<$1.052 is determined, resulting in work extraction $>$91% of perfect isothermal work. Thermodynamic simulations indicate that the proposed heat engine produces up to 22.6% greater power output than an ORC-based counterpart for a 100–374 ${}^{\circ}C$ constant temperature heat source. Importantly, the isothermal expansion of the cycle facilitates the use of ’wet’ fluids such as water—fluids that challenge the performance in ORC systems but are well-matched for high-temperature heat sources. Since the turbine’s practical efficiency is obscure, we demonstrate that turbine efficiencies between 70% and 90% yield an improvement of 42.7% and 90.7%, respectively, compared to the ORC heat engines with turbine isentropic efficiency of 75% at the temperature of 373 ${}^{\circ}C$. These findings provide a basis for developing efficient heat engines tailored to heat-to-power applications.
RESULTS AND DISCUSSION
Nozzle design
The nozzle is designed according to a previously developed model for two-phase flow in a converging-diverging nozzle 21. The design is based on the homogenous model of a bubbly mixture flowing in the nozzle, which was shown valid by others 22, 23, 24, 17, 18. Under this model, we assume that the phases are in pressure, temperature, and velocity equilibrium. The geometry of the cross-section is selected as circular. In contrast to the propulsion nozzle, where the pressure at the injection chamber is dictated and limited by the cruise velocity, the inlet pressure is a free parameter controlled by a pump, which results in supersonic flow. To illustrate the physical principle and evaluate the expansion, we introduce air into a water stream at ambient temperature. The injection pressure at the design point was chosen to be 6 bar, an air mass flow rate of 13.1 $g/s$, an injection temperature of 26 ${}^{\circ}C$, and a water volumetric flow rate of 14.6 $m^{3}/hr$. The nozzle is designed to implement a constant $dp/dt$ to ease the heat transfer between the phases. Figure 1 shows a schematic representation of the designed nozzle.
Figure 1. Nozzle design drawing
Near-isothermal expansion
During the experiments, static pressure was measured at the nozzle’s injection, outlet, and other locations along the nozzle. Figure 2 presents the pressure distribution and the nozzle’s diameter in the nozzle (solid black curve). It indicates that supersonic conditions are met in the nozzle since the pressure continues to drop after the throat until it reaches atmospheric pressure at the nozzle’s outlet. This pressure distribution matches the nozzle’s design, i.e., reducing the pressure from a 6 bar in the air injection chamber to atmospheric pressure in the nozzle’s outlet. Since isothermal expansion was assumed in the nozzle’s design, this is the first reinforcement of our isothermal process assumption. We note that the pressure measurement at the throat stands out, where the measured pressure is considerably lower than anticipated. We ascribe this discrepancy to the high throat sensitivity to manufacturing flaws, slip 25, and tap diameter 26 due to high pressure and velocity gradients in this region. A computational fluid dynamics (CFD) simulation was performed to analyze the anticipated pressure distribution at the working conditions, which is detailed in the experimental procedure subsection. The CFD outcomes closely align with the experimental data, corroborating the isothermal expansion hypothesis, as the CFD model assumes uniform mixture temperature.
Figure 2. Pressure distribution comparison between the measured, simulated, and designed values
Subsequently, we employ the polytropic index to gauge the extent to which this process approximates a strictly isothermal expansion. To quantify the polytropic index, we apply the energy conservation equation to the nozzle’s outlet, where the pressure is P, and the mixture’s velocity is expressed by:
$u^{2}=\frac{2\dot{V}_{w}(P_{inj}-P)}{\dot{m}_{w}+\dot{m}_{a}}+u_{inj}^{2}+\frac{2}{\dot{m}_{w}+\dot{m}_{a}}\times(\dot{W}_{a}-\dot{E}_{loss})$ (1)
where $\dot{V}_{w},P_{inj},\dot{m}_{a},\dot{m}_{w},u_{inj},\dot{W}_{a}$ and $\dot{E}_{loss}$ are the water volumetric flow rate, air injection pressure, air mass flow rate, water mass flow rate, water injection velocity, work rate exerted by air, and energy losses, respectively. Further discussion on the governing equations can be found in the experimental procedure section. The air expansion work is assumed to be a polytropic process, for which the work rate is calculated by 12:
$\dot{W}_{a}=P_{inj}\dot{V}_{a,inj}\frac{1-(\frac{P_{inj}}{P_{out}})^{\frac{1-n}{n}}}{n-1}$ (2)
where n and $P_{out}$ are the polytropic index and the outlet pressure, respectively. As a result, equation 1 involves two unknown variables: the polytropic index and the head losses in the nozzle.
To gain insights on the head losses in the nozzle, we evaluate the expression $\dot{E}_{loss}=\dot{V}_{w}\Delta P$, where $\Delta P$ is the pressure loss, and is calculated by $\Delta P=\zeta\frac{\rho u^{2}}{2}$. Here, $\zeta$, $\rho$, and $u$ are the resistance coefficient, mixture’s density and inlet velocity, respectively 27.
In the above equation, $\zeta$ depends on geometrical factors while the term $\rho u^{2}$ strongly depends on the void fraction and, thus, the temperature and the polytropic index. Consequently, comprehending the influence of the polytropic index on this parameter is sufficient to infer the nuzzle’s head loss characteristics. This relationship is quantified by the equation below:
$\rho u^{2}=\left[\alpha\rho_{a}+(1-\alpha)\rho_{w}\right]\times\frac{\dot{V}^{2}_{w}}{A^{2}(1-\alpha)^{2}}=\frac{\dot{V}^{2}_{w}}{A^{2}}\times\left[\frac{\alpha\rho_{a}}{(1-\alpha)^{2}}+\frac{\rho_{w}}{1-\alpha}\right],$ (3)
where $\alpha$ is the void fraction and $A$ is the cross-sectional area. Figure 3 plots the above term as a function of the polytropic index using the measured conditions at the nozzle’s outlet. The head losses monotonically reduce with increasing the polytropic index, indicating that the largest head losses are received for the pure isothermal case. Therefore, the head losses obtained from the CFD simulation, representing perfect heat transfer between the phases, can be considered an upper bound for the actual head loss.
Figure 3. Head loss dependence on the polytropic index
In Figure 4, we plot the solution for the energy conservation equation (equation 1) for the nominal conditions and water volumetric flow rates error of 0.5%. Since both head losses and $n$ are unknown, each point on this curve is a solution to this equation. Using equation 1, the head losses derived from the CFD simulation is 0.84 bar. This value is used as an upper bound to our experiment; thereby, we can extract an upper bound on the polytropic index, $n$. From this analysis, the upper bound for the polytropic index is 1.052.
Figure 4. Solution to equation 1 using measured data.
One of the benefits of isothermal expansion stems from the additional work that can be extracted in the expansion process. Hence, It is instructing to understand the improved extracted work compared to the adiabatic process, which represents gas expansion with no heat exchange. Therefore, we define the isothermal efficiency by the ratio between the polytropic-index dependent work, $\dot{W}_{a}=P_{inj}\dot{V}_{a,inj}\frac{1-(\frac{P_{inj}}{P_{out}})^{\frac{1-n}{n}}}{n-1}$, and the isothermal work $\dot{W}_{iso}=P_{inj}\dot{V}_{a,inj}ln\left(\frac{P_{inj}}{P_{out}}\right)$ 12:
$\eta_{iso}=\frac{\dot{W}_{a}}{\dot{W}_{iso}}=\frac{1-r_{p}^{\frac{1-n}{n}}}{(n-1)\mathrm{ln}(r_{p})},$ (4)
where $P_{out}$ is the outlet pressure and $r_{p}=\frac{P_{inj}}{P_{out}}$ is the pressure ratio.
Figure 5 shows the isothermal efficiency using the extracted polytropic index of $n=1.052$ for pressure ratios up to 10, for which isothermal efficiencies $>94\%$ are attained. As such, an additional 44-71 % more work can be extracted compared to adiabatic expansion, for which $n=1.4$. In the context of heat engines using organic working fluids, these are associated with a lower polytropic index of $n=1.05-1.3$. As such, increasing work extraction by approaching isothermal expansion is limited and dependent on the polytropic index of the working fluid.
Figure 5. Isothermal efficiency for varying pressure ratios.
High impeller efficiency
The near-isothermal expansion of bubbles experienced in our nozzle reveals opportunities for a new type of heat engine utilizing such nozzles for a reaction turbine. The turbine consists of an impeller connected to nozzles. The rotor leg is submerged in the HTL reservoir, sucking HTL into the rotor. Similarly to a centrifugal pump impeller, the impeller vanes pressurize the HTL until reaching maximal static pressure at the end of the impeller, where the nozzles begin, and liquid WF is injected and mixed with the HTL. However, unlike the centrifugal pump, our impeller has no stator, which minimizes energy losses while maintaining high design flexibility. In the nozzles, heat is transferred from the HTL to the WF until the WF is fully evaporated, and the mixture is accelerated until it reaches the nozzles’ outlet, where the jet flows tangentially, creating a thrust force and rotates the shaft for electricity generation. An example of the impeller design is shown in Figure 6.
Figure 6. Impeller design. (a) hidden shroud, and (b) with shroud.
Figure 7 exemplifies CFD simulation of an impeller design for 10.5 bar injection pressure, resulting in an impeller efficiency of $\sim 97\%$.
Figure 7. CFD results of the impeller at 1500 RPM. Pressure values are expressed in bar gauge.
Description of the ISO cycle
This subsection details the novel thermodynamic cycle (ISO cycle henceforth). In this cycle, depicted in Figure 8, the HTL transfers thermal energy from an outside heat source to the WF. Thermal energy is converted into kinetic energy at the nozzle, creating thrust and rotating a reaction turbine to generate electricity.
The heat source passes through a heat exchanger (1$\rightarrow$2), which transfers heat to the HTL (3$\rightarrow$4). The HTL is pumped by the turbine in conjunction with colder HTL from the reservoir (5). The HTL flow from the reservoir is used to obtain the desired HTL flow rate in the nozzle, resulting in a specific void fraction. Before reaching the nozzle, the impeller increases the HTL’s pressure to the WF’s saturation pressure. The HTL reaches the nozzle’s inlet and mixes with colder WF liquid (6) in the nozzle’s injection chamber. Heat transfers to the WF, which raises its temperature to its evaporation temperature. The WF then completely evaporates and expands inside the nozzle within the HTL until it reaches its outlet (7). This expansion raises the mixture’s velocity and generates thrust, which turns the reaction turbine. The rotor rotates in a tank that functions as a vertical vapor-liquid separator, separating the WF from the HTL after the expansion by generating a film flow of the mixture on the tank boundaries. The hot WF gas enters a recuperator, which transfers heat from the hot WF to the cold condensed WF before it enters the condenser (8$\rightarrow$9). This internal regeneration, facilitated by the isothermal expansion, is used to heat the cold liquid WF, minimizing heat transfer irreversibilities. Then, the liquid WF is pumped through the recuperator (10$\rightarrow$11). The WF is then injected into the nozzles, mixing with the hot HTL (6).
Figure 8. Schematic of the ISO cycle.
ISO cycle performance analysis
In the ISO cycle, isothermal expansion drives vapors away from saturation conditions, ensuring the fluid remains in a single phase. This, in turn, enables the use of ’wet’ fluids—like water—that have a higher critical temperature than the ’dry’ fluids. This lies in contrast to ORC-based heat engines, where it is advisable to employ ’dry’ fluids to limit superheat and sustain a dry turbine 28.
Therefore, the ISO cycle was simulated with water as the WF and a heat source in 100-374 C, covering a substantial portion of the waste heat recovery temperature range. The simulated heat source is characterized by a constant temperature, representing condensation scenarios harnessing constant-temperature latent heat. This application is optimally suited for the ISO cycle, as the predominant portion of the heat involved is latent heat, allowing significant heat extraction with minimal temperature change. The scale of the system is a few hundred kW, which is the scale needed for many industrial waste heat recovery systems. The heat sink temperature was chosen as 20 C, typical for atmospheric conditions.
To understand the proposed cycle’s benefits, we compare its performance to a conventional recuperated ORC. In the simulated ORC cycle, the WF is heated in a boiler by the same heat source and expands in a turbine. Then, the colder WF transfers heat by a recuperator before it condenses and is pumped back into the boiler through the recuperator. In the two cycles, the temperature of the hot WF is optimized to extract maximum power.
The ORC is simulated with cyclopentane as the WF using the same heat source. Cyclopentane is chosen due to its high critical and working temperatures of 238.54C and 300C, respectively 29. Furthermore, cyclopentane is suggested as a dry fluid for the organic Rankine cycle for low-grade heat conversion 28. The thermophysical properties of the materials were acquired from the National Institute of Standards and Technology (NIST) REFPROP database 30.
Figure 9 exemplifies T-S diagrams of the two cycles using heat source temperature ($T_{H}$) of 200 C with cyclopentane (a) and water (b). The larger ISO cycle in Figure 9 a implies that more output power is available than the conventional ORC per WF unit mass flow. The isothermal expansion allows a more significant internal regeneration using a recuperator (289 VS 39 kJ/kg), and as such, the liquid preheating portion is reduced, and more heat is transferred isothermally to the WF. Consequently, heat transfer irreversibilities are reduced compared to the ORC, resulting in increased thermal efficiency. Additionally, no superheating is needed in the ISO case, further reducing irreversibilities and increasing the turbine’s pressure ratio.
Figure 9. T-S diagrams of the compared cycles at $T_{H}=200^{\circ}C$. a) Cyclopentane comparison. b) ISO cycle with water as the WF.
[…]
Figure 10. Comparative analysis between the ISO cycle and ORC of efficiencies (a) and net output work (b).
Table 1. ISO cycle (water) improvement
$T_{H}$ Power output improvement Specific work improvement
(C) (%) (%)
100 12.3 591
120 14.0 609
140 15.5 627
160 17.0 645
180 18.6 665
200 20.5 690
220 21.6 707
240 21.0 708
260 21.1 717
280 21.5 727
300 21.8 738
320 22.1 749
340 22.3 759
360 22.6 769
373 22.6 773
Since the reaction turbine has yet to be built and experimented with, we consider turbine efficiencies in the range of 20-90%, where the ISO turbine efficiency is defined by the ratio of the turbine output power to the WF expansion work. In this context, we assume the entropy produced due to turbine irreversibilities is retained within the system in heat. In contrast to the ORC turbine, where the generated entropy cannot be utilized, the ISO cycle incorporates this thermal energy into the HTL stream, thereby augmenting the flow rate of the working fluid. This process mitigates the decline in work output due to turbine inefficiencies. Figure 11 illustrates the output power enhancement of the ISO cycle, using water as the working fluid, across the varying turbine efficiencies relative to the ORC, considering a realistic ORC expander exhibits isentropic efficiency of 75%, consistent with values used in past studies 4. The black curve in the figure denotes the ISO turbine efficiency for which the output power of the two cycles is equal. The figure indicates that in the considered ISO turbine efficiencies, the ISO cycle potentially generates up to 90.7% more power than the ORC-based heat engine and is superior to the ORC-based heat engines for ISO turbine efficiencies exceeding 47% in the examined temperature range.
Figure 11. ISO cycle power output improvement compared to the ORC as a function of the heat source temperature and the ISO turbine efficiency considering ORC turbine isentropic efficiency of 75%. The black curve denotes the ISO turbine efficiency for which the ISO cycle and ORC produce equal power.
When applied to ideal gas cycles, such as the Ericsson cycle, the proposed turbine may offer even higher-efficiency heat engines. There is no phase change in this cycle, and the gas undergoes isothermal compression, theoretically reaching Carnot’s efficiency. Thus, the successful implementation of effective isothermal compression could pave the way for developing heat engines with efficiencies surpassing those achievable by the ISO cycle.
Conclusions
This study introduces the concept of a supersonic two-phase nozzle as a key component of a novel thermodynamic cycle utilizing isothermal expansion. Experimental pressure measurements with air in water validated the supersonic conditions, while CFD allowed us to extract an upper bound for the polytropic index of air of $n<1.052$. This index indicates quasi-isothermal expansion, enabling up to 71% more work extraction than adiabatic expansion. Compared to other reported methods, such as flooded expansion, this method achieves a much higher liquid-to-working-fluid mass ratio (0.997 compared to 0.3). This significantly increases the heat transfer between the phases, enabling almost isothermal gas expansion. Furthermore, using the two-phase nozzle is an important step for increasing the pressure ratio, which makes it more suitable for various applications.
To leverage this nozzle for realizing a heat engine, we theoretically explored a reaction turbine, which rotates due to thrust created by the nozzles. CFD analysis of the impeller shows hydraulic efficiency of $\sim 97\%$, supporting its efficiency potential compared to existing small-scale turbines.
The proposed heat engine, based on the novel cycle (termed ISO cycle), yielded superior thermodynamic performance compared to a heat engine based on a conventional ORC cycle and showed up to 22.6% higher power output for heat source temperatures up to 373C.
We broadened our simulation to encompass turbine efficiencies spanning from 20% to 90% and compared to ORC small-scale heat engine, which encompasses turbine isentropic efficiency of 75%. This analysis revealed that the ISO cycle outperforms the ORC within the examined temperature range when turbine efficiency exceeds 47%.
Not only is it more efficient, but the ISO turbine has the potential to be much smaller than an ORC counterpart. This is because the energy carrier in the ISO cycle is liquid, three orders of magnitude denser than gas, which is the typical energy carrier in ORC and other heat engines.
Future research should focus on building and experimentally validating the proposed cycle, optimizing its components, and exploring its potential applications in various energy systems.
Experimental procedures
Nozzle CFD analysis
A 3D CFD simulation using the finite volume methods was performed on a nozzle using the commercial solver FLUENT 2024R1. The CFD model was validated by comparing the static pressure results with the experimental measurements along the nozzle. Due to the highly dispersed gas phase in the liquid flow and the high boundary area between phases, a ’Mixture’ multiphase model was used. The ’Mixture’ multiphase model is a simplified model that treats the mixture as a single continuum without sharp boundaries between the phases. The standard fluid dynamic governing equations are modified by introducing the void fraction parameter. No slip velocity between phases was assumed. A steady Reynolds Averaged Navier-Stokes (RANS) turbulence modeling approach was used to minimize simulation runtime. The RANS approach provides a solution for the time-averaged turbulent flow parameters. Although it eliminates the fluctuating terms, it is a powerful simplification that allows reasonably accurate results under reasonable computation costs. As a closure model, K-omega SST was compared with the K-epsilon models. Both models produced similar results. The simulations were performed with several cell sizes to verify the independence of the results from cell size. Polyhidara cells were used. Prism cells were modeled for the wall treatment. Using the Effective Viscosity Ratio (EVR) parameter, a sufficient number of prism cells was assured, ensuring the actual boundary layer falls within the scope of the prism cells.
Impeller CFD analysis
The impeller is designed and optimized using ANSYS 2024R1 software. For this purpose, a few ANSYS modules were used. We used Vista CPD to define the initial impeller geometry for specific operating conditions, followed by BladeGen to create the final blade, shroud, and hub geometries. In the next step, the fast solver of VistaTF is applied to check operating conditions. Finally, Fluent’s full CFD solution is done to evaluate impeller efficiency and $NPSH_{r}$. The fine mesh with sufficient boundary layer modeling was used, and three levels of mesh refinement validated the convergence. For all meshes, y+ was less than 1.
ISO cycle Simulation model
In this subsection, we outline the thermodynamic framework applied to assess the ISO cycle.
For each temperature, we optimized the temperature $t_{3}$ for maximum output power and solved the following equations:
  1. 1.
    $\dot{w}_{exp}=t_{7}(s_{7}-s_{6})-(h_{7}-h_{6})$
  2. 2.
    $\dot{q}_{in}=h_{6}-h_{11}+t_{8}(s_{7}-s_{6})$
  3. 3.
    $\dot{m}_{WF}=\frac{\dot{Q}_{in}}{\dot{q}_{in}}$
where $s$, $h$, $t$, $\dot{q}$, $\dot{w}$, $\dot{V}$ and $\dot{m}$ are the specific entropy, enthalpy, temperature, heat, work, volumetric flow rate and mass flow rate, respectively, and the numbered subscripts denote the thermodynamic cycle points presented in Figure 8. $\dot{Q}_{in}$ is the rate heat transferred into the system.
For each heat source temperature, we optimize the nozzle’s outlet temperature to achieve maximum net power output.
The thermal efficiency of the cycle is calculated as follows:
$\eta_{th}=\frac{\dot{W}_{net}}{\dot{Q}_{in}},$ (5)
The net output work of the heat engine is calculated as:
$\dot{W}_{net}=\dot{W}_{T}-\dot{W}_{p,WF}$ (6)
where $\dot{W}_{T}$ and $\dot{W}_{p,WF}$ are the turbine’s output work and the work consumed by the WF pump, respectively.
The turbine’s output work is calculated as:
$\dot{W}_{T}=\dot{W}_{exp}\times\eta_{T,ISO}$ (7)
where $\dot{W}_{exp}$ and $\eta_{T,ISO}$ are the expansion work and ISO turbine efficiency, respectively.
RESOURCE AVAILABILITY
Lead contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Dror Miron (drormiron@gmail.com).
Materials availability
This study did not generate new materials.
Data and code availability
  • Any information required to reanalyze the data reported in this paper is available from the lead contact upon request.
ACKNOWLEDGMENTS
The work of D. Miron was supported by the Israel Department of Energy (PhD fellowship) and by the Nancy and Stephen Grand Technion Energy Program (GTEP). The authors acknowledge the contribution of Matan Hameiry to this work.
AUTHOR CONTRIBUTIONS
Conceptualization, D.M., Y.N, J.C., and C.R; methodology, D.M., Y.N, J.C., and C.R.; investigation, D.M., Y.N, N.F, A.S, and J.C.; visualization, D.M.; writing-–original draft, D.M., Y.N, J.C., and C.R.; writing-–review & editing, D.M., Y.N, J.C., A.S, N.F, and C.R.; funding acquisition, C.R.; supervision, C.R.
DECLARATION OF INTERESTS
All authors are employees and shareholders of Lava Energy Ltd. Carmel Rotschild is a founder of Lava Energy and a member of its scientific advisory board. The authors have submitted the following relevant patents:
  • HEAT ENGINE WO2022/049573
  • TWO PHASE HEAT ENGINE WO2023228173
  • Heat engine using a liquid-vapor-phase-changing material WO2024/028878